# TabDistill: Distilling Transformers into Neural Nets for Few-Shot Tabular Classification

### v1

@@ 1 Introduction
Toward answering this question, we propose TabDistill, a novel framework to distill the classification capabilities of pre-trained transformers into neural networks for few-shot tabular classification (also see Figure 1). We draw inspiration from the image domain where transformer-based models have been found to be good hypernetworks for generating neural networks to implicitly represent images [Chen2022TransINR, gu2025enhancingFoundationModelsHypernetworks]. Succinctly, TabDistill incorporates the pre-trained knowledge of a transformer-based model (the base model) into a neural network by fine-tuning the transformer to infer its weights. We assume that the base model contains an informative intermediate representation (for example, the encoder output of TabPFN [hollmann2023TabPFN] or encoder-decoder type language models such as BERT [devlin2019bert], BART [lewis2019bart], BigScience T0 series [sanh2021t0Models] etc.). TabDistill learns a linear map for projecting the intermediate representation provided by the base model into the parameter space of the neural network, by fine-tuning using the cross-entropy loss of the resultant classifier, and achieves both data efficiency and parameter efficiency. We employ a novel permutation-based training technique to avoid overfitting the model to the extremely small number of training examples.
@@ 1.1 Related Works
Meta-Learning and hypernetworks. Meta-learning is the process of learning to generalize to unseen tasks by observing few examples corresponding to each task [vilalta2002metaLearningPerspective]. Transformers are known to be good at meta learning, particularly in the form of in-context learning [kirsch2022metaLearningTransformers]. Hypernetworks are closely related meta-learning, in the sense that they predict parameters for other machine learning models by observing a few samples from the task at hand. Transformers have been used as hypernetworks in computer vision applications, specifically for generating implicit neural representations [Chen2022TransINR, gu2025enhancingFoundationModelsHypernetworks]. In [Chen2022TransINR], a transformer is trained from scratch to predict weights of a neural network which represents an image or a 3D scene. Another work [gu2025enhancingFoundationModelsHypernetworks] exploits the pre-trained knowledge of a transformer-based foundation model for the same task. Both the works append additional placeholder tokens to the input for predicting the neural network weights. In contrast, our framework directly maps the embedding space to neural network parameters.
Knowledge distillation. Knowledge distillation [gouKDSurvey, sucholutsky2023getting, hintonDistillation, romeroFitnets, dissanayakequantifying, xu2024survey, liang2023less, hamman2025few] is another popular paradigm which trains a smaller student model from a larger teacher model (often from the same model family) by directly aligning the student’s outputs with the teacher’s outputs using a loss function. Our setup is distinctly different since we extract a neural network out of a transformer-based model without any loss-function-based alignment.
@@ 2.1 Notation and problem setup
The MLP $h_{\theta}(x)$ has the following architecture. Let $\texttt{ReLU}(u)$ denote the ReLU activation function [glorot2011relu]. With the hyperparameters $R$ and $L$ denoting the number of layers and the width of the hidden layers, respectively, $h_{\theta}(x)$ is defined as
@@ 2.2 Our proposed TabDistill framework
    Create a combined serialized input $g(\mathcal{D}_{N})$ for the complex model $f$ using the training data $\mathcal{D}_{N}$
@@
    Prompt the complex model encoder to get the embeddings $z=f_{E}(g(\mathcal{D}_{N}))$
@@
    Infer the parameters $\theta=m_{\eta}(z)$ for the MLP $h_{\theta}$
@@
    Compute the loss $\mathcal{L}(\eta;\mathcal{D}_{N})$ as the cross entropy loss of the classifier $h_{\theta}$
Ultimately, the fine-tuning loss function can be written as

### v2

@@ 1 Introduction
Toward answering this question, we propose TabDistill, a novel framework to distill the classification capabilities of pre-trained transformers into neural networks for few-shot tabular classification (also see Figure 1). We draw inspiration from the image domain where transformer-based models have been found to be good hypernetworks for generating neural networks to implicitly represent images [Chen2022TransINR, gu2025enhancingFoundationModelsHypernetworks]. Succinctly, TabDistill incorporates the pre-trained knowledge of a transformer-based model (the base model) into a neural network by fine-tuning the transformer to infer its weights. We assume that the base model contains an informative intermediate representation (for example, the encoder output of TabPFN [hollmann2023TabPFN] or BERT [devlin2019bert], encoder-decoder type language models such as BART [lewis2019bart] and BigScience T0 series [sanh2021t0Models] etc.). TabDistill learns a linear map for projecting the intermediate representation provided by the base model into the parameter space of the neural network, by fine-tuning using the cross-entropy loss of the resultant classifier, and achieves both data efficiency and parameter efficiency. We employ a novel permutation-based training technique to avoid overfitting the model to the extremely small number of training examples.
@@ 1.1 Related Works
Meta-Learning and hypernetworks. Meta-learning is the process of learning to generalize to unseen tasks by observing few examples corresponding to each task [vilalta2002metaLearningPerspective]. Transformers are known to be good at meta learning, particularly in the form of in-context learning [kirsch2022metaLearningTransformers]. Hypernetworks are closely related meta-learning, in the sense that they predict parameters for other machine learning models by observing a few samples from the task at hand. Transformers have been used as hypernetworks in computer vision applications, specifically for generating implicit neural representations [Chen2022TransINR, gu2025enhancingFoundationModelsHypernetworks]. In [Chen2022TransINR], a transformer is trained from scratch to predict weights of a neural network which represents an image or a 3D scene. Another work [gu2025enhancingFoundationModelsHypernetworks] exploits the pre-trained knowledge of a transformer-based foundation model for the same task. Both the works append additional placeholder tokens to the input for predicting the neural network weights. In contrast, our framework directly maps the embedding space to neural network parameters. MotherNet [mueller2025mothernet] is a hyper-network based on a TabPFN-like model which provides a neural network classifier in a single forward pass with in-context learning, given a set of tabular training data. It is pre-trained on a large number of synthetic tabular data. After the initial pre-training phase (i.e., the meta-learning phase), MotherNet does not need any additional training to adapt to an unseen dataset. In contrast to our work, the focus is not on the few-shot regime, and the architecture of the generated neural network is fixed irrespective of the actual dataset and its features.
Knowledge distillation. Knowledge distillation [gouKDSurvey, sucholutsky2023getting, hintonDistillation, romeroFitnets, dissanayakequantifying, xu2024survey, liang2023less, hamman2025few] is another popular paradigm that trains a smaller student model from a larger teacher model (often from the same model family) by directly aligning the student’s outputs with the teacher’s outputs using a loss function. Our setup is distinctly different since we extract a neural network out of a transformer-based model without any loss-function-based alignment.
@@ 2.1 Notation and problem setup
The MLP $h_{\theta}(x)$ has the following architecture. Let $\texttt{ReLU}(u)$ denote the ReLU activation function [glorot2011relu]. With the hyperparameters $R$ and $L$ denoting the number of layers and the width of the hidden layers, respectively, $h_{\theta}(x)$ is defined as:
@@ 2.2 Our proposed TabDistill framework
    Create a combined serialized input $g(\mathcal{D}_{N})$ for the complex model $f$ using the training data $\mathcal{D}_{N}$.
@@
    Prompt the complex model encoder to get the embeddings $z=f_{E}(g(\mathcal{D}_{N}))$.
@@
    Infer the parameters $\theta=m_{\eta}(z)$ for the MLP $h_{\theta}$.
@@
    Compute the loss $\mathcal{L}(\eta;\mathcal{D}_{N})$ as the cross entropy loss of the classifier $h_{\theta}$.
Ultimately, the fine-tuning loss function can be written as: